OpenStudy (anonymous):

In a circle centered at point O, the ratio of the area of sector AOB to the area of the circle is 3/5. What is the approximate measure, in radians, of the central angle corresponding to AB? Round the answer to two decimal places.

OpenStudy (anonymous):

so basically, pi*r^2* x/360 = 3/5* pi*r^2 so x/360 = 3/5 when you cancel things so x = 360*3/5 = 216 i think 3.77 radians

OpenStudy (anonymous):

thank you!!! and thanks for writing out the formula!

OpenStudy (anonymous):

haha welcome :) all tthe best :)

OpenStudy (anonymous):

I understand how you got the 216 just not the 3.77